You are long 10 gold futures contracts, established at an initial settle price of $1,620 per ounce, where each contract represents 100 ounces. Over the subsequent four trading days, gold settles at $1,626, $1,619, $1,623, and $1,626, respectively.

  

a.

Calculate the profit or loss for each trading day. (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.)

b.

Compute your total profit or loss at the end of the trading period. (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.)
Answer #1

a). Initial Value = No. of Contracts * Total ounces in each contract * Settle Price

= 10 * 100 * $1,620 = $1,620,000

Day 1 account value = 10 * 100 * $1,626 = $1,626,000

Day 1 Profit/Loss = Day 1 account value - Initial Value = $1,626,000 - $1,620,000 = $6,000

Day 2 account value = 10 * 100 * $1,619 = $1,619,000

Day 2 Profit/Loss = Day 2 account value - Day 1 account value = $1,619,000 - $1,626,000 = -$7,000

Day 3 account value = 10 * 100 * $1,623 = $1,623,000

Day 3 Profit/Loss = Day 3 account value - Day 2 account value = $1,623,000 - $1,619,000 = $4,000

Day 4 account value = 10 * 100 * $1,626 = $1,626,000

Day 4 Profit/Loss = Day 4 account value - Day 3 account value = $1,626,000 - $1,623,000 = $3,000

b). Total Profit/Loss = Day 4 account value - Initial value = $1,626,000 - $1,620,000 = $6,000